# Beam Enumeration: Probabilistic Explainability For Sample Efficient Self-conditioned Molecular Design

### v1

@@ Appendix A Beam Enumeration
A.1 Algorithm Overview
Figure A4: Beam Enumeration overview. a. The proposed method proceeds via 4 steps: 1. generate batch of molecules. 2. filter molecules based on pool to enforce substructure presence, discarding the rest. 3. compute reward 4. update the model. After updating the model, if the reward has improved for consecutive epochs, execute Beam Enumeration. b. Beam Enumeration sequentially enumerates the top $k$ tokens by probability for $N$ beam steps, resulting in an exhaustive set of token sub-sequences. c. All valid substructures (either by the Structure or Scaffold criterion) are extracted from the sub-sequences. The most frequent substructures are used for self-conditioned generation. This overview figure is the same as in the main text.
Beam Enumeration (Fig. A4) is an algorithm that extracts molecular substructures from a generative model’s weights for self-conditioned generation. The problem set-up is any molecular design task to optimize for a target property profile, e.g., high predicted solubility and binding affinity. When molecular generative models are coupled with an optimization algorithm, it should be increasingly likely to generate desirable molecules, i.e., molecules that possess the target property profile.
@@ A.1 Algorithm Overview
In this work, Beam Enumeration is applied to a language-based autoregressive generative model operating on the simplified molecular-input line-entry system (SMILES)Weininger (1988) representation. The optimization algorithm is Augmented MemoryGuo and Schwaller (2023) which builds on REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a) and casts the optimization process as an on-policy reinforcement learning (RL) problem. Following RL terminology, sampling from the generative model involves sampling trajectories, which in this case, are SMILES, and the desirability of the corresponding molecule is given by the reward.
The underlying hypothesis of Beam Enumeration is that during the RL optimization process, partial trajectories provide a source of signal that can be exploited. Usually, full trajectories are sampled which map to a complete SMILES sequence that can be translated to a molecule. Our assumption is that partial trajectories (partial SMILES sequence) can be mapped to molecular substructures (a part of the full molecule). This statement is not guaranteed as SMILES and molecules are discrete and small perturbations often leads to invalid SMILES. We prove this assumption in Section C by showing that although the vast number of partial trajectories do not map to valid SMILES, the raw number is sufficient to extract a meaningful signal. Correspondingly, Beam Enumeration leverages partial trajectories on the assumption that molecular substructures are on track to becoming full molecules that would receive high reward.
A.2 Enumerating Partial Trajectories
@@
A.3 Extracting Molecular Substructures
Given a set of token sub-sequences, the goal is to extract out the most frequent molecular substructures. This is done by taking each sequence, considering every (sub)-sub-sequence, and counting the number of valid substructures (Fig. A). For example, given the sub-sequence “ABCD”, the set of (sub)-sub-sequences are: “A”, “AB”, “ABC”, and “ABCD”. In practice, we only consider (sub)-sub-sequences with at least three characters (“ABC” and “ABCD”) since each character loosely maps to one atom and three is approximately the minimum for meaningful functional groups, e.g., “C=O”, a carbonyl. The set of most frequent substructures is assumed to be on track to receive a high reward.
A.4 Defining Molecular Substructures: Scaffold vs. Structure
As shown in Fig. A, molecular substructures can be defined on the Scaffold or Structure level. The former extracts the Bemis-MurckoBemis and Murcko (1996) scaffold while the latter extracts any valid structure. The any valid structure is an important distinction as our experiments find that extracting by Structure leads to the most frequent molecular substructures being small functional groups that do not have corresponding scaffolds. By contrast, extracting the scaffold always leads to ring structures. Moreover, extracting specifically the Bemis-Murcko scaffold is important as heavy atoms, e.g., nitrogen, are important for biological activity. Consequently, extracted substructures are also enforced to contain at least one heavy atom as we find that benzene, perhaps unsurprisingly, is commonly the most frequent substructure. See Section B for more details on the differing behavior of ‘Scaffold’ vs. ‘Structure’.
A.5 Self-conditioned Generation
@@
A.6 Probabilistic Explainability
The set of most frequent molecular substructures should be meaningful as otherwise, the model’s weights would not have been updated such that these substructures have become increasingly likely to be generated. We verify this statement in the illustrative experiment in the main text and in Section C. In the drug discovery case studies (Appendix D), the extracted substructures are more subtle in why they satisfy the target objective but certainly must possess meaning, however subtle, as otherwise, they would not receive a high reward. In the main text, we show that extracted substructures form core scaffolds and structural motifs in the generated molecules that complement the protein binding cavity. Finally, we emphasize that the correctness and usefulness of this explainability deeply depends on the oracle(s) being optimized for. The extracted substructures do not explain why the generated molecules satisfy the target objective. Rather, they explain why the generated molecules satisfy the oracle. The assumption in a generative design task is that optimizing the oracle is a good proxy for the target objective, e.g., generating molecules that dock well increases the likelihood of the molecules being true binders. This observation directly provides additional commentary on why sample efficiency is so important: the ability to directly optimize expensive high-fidelity oracles would inherently enhance the correctness of the extracted substructures.
Appendix B Beam Enumeration: Findings from Hyperparameter Screening
@@
B.1 Beam Enumeration Hyperparameters
@@
Substructure Type. This hyperparameter has two possible values: Scaffold or Structure. Scaffold extracts Bemis-MurckoBemis and Murcko (1996) scaffolds while Structure extracts any valid substructure.
Structure Structure Minimum Size. This hyperparameter enforces the partial SMILES to contain at least a certain number of characters. In effect, this enforces extracted molecular substructures to be larger than a Structure Minimum Size. From the illustrative experiment in the main text and Section C, Structure extraction often leads to small functional groups being the most frequent in the sub-sequences. By enforcing a minimum structure size, Structure extraction leads to partial structures which may carry more meaning. We find that this hyperparameter greatly impacts sample efficiency and we present all our findings in a later sub-section.
@@
Patience. This hyperparameter controls how many successive reward improvements are required before Beam Enumeration executes and molecular substructures are extracted. Recalling the first fact in which Beam Enumeration was proposed on: On a successful optimization trajectory, the model’s weights must change such that desirable molecules are more likely to be generated, on average. Patience is effectively an answer to "when would extracted substructures be meaningful?" Too low a patience and stochasticity can lead to negative effects while too high a patience diminishes the benefits of Beam Enumeration on sample efficiency.
Token Sampling Method. This hyperparameter has two possible values: "topk" or "sample" and denotes how tokens sub-sequences are enumerated. "topk" takes the top k most probable tokens at each Beam Step while "sample" samples from the distribution just like during batch generation. Our results show interesting observations surrounding this hyperparameter as "sample" can work just as well and sometimes even better than taking the "topk". These results were unexpected as the underlying hypothesis is that the most probable sub-sequences lead to the most useful substructures being extracted. However, our findings are not in contradiction as sampling the conditional probability distributions would still lead to sampling the top k tokens, on average. Moreover, after extensive experiments, we find that "sample" leads to more variance in performance across replicates which is in agreement with the assumption that sampling the distributions can lead to more improbable structures. We thoroughly discuss our findings in a later sub-section where we provide hyperparameter recommendations and analyses to the effects of tuning each hyperparameter.
B.2 Hyperparameters: Grid Search
@@
Figure B5: illustrative experiment Generative Yield > 0.8. The IntDiv1Polykovskiy et al. (2020a) is annotated.
Figure B6: illustrative experiment Generative Yield > 0.8 with Structure Minimum Size. The IntDiv1Polykovskiy et al. (2020a) is annotated.
Figure B7: illustrative experiment Generative Yield > 0.8 with Structure Minimum Size and "Sample" token sampling. The IntDiv1Polykovskiy et al. (2020a) is annotated.
Figure B8: illustrative experiment Oracle Burden (10) > 0.8
Figure B9: illustrative experiment Oracle Burden (10) > 0.8 with Structure Minimum Size
Figure B10: illustrative experiment Oracle Burden (10) > 0.8 with Structure Minimum Size and "Sample" token sampling
B.3 Analysis of Grid Search Results
@@
  • Scaffold and Structure with Structure Minimum Size: "sample" sampling can be better than "topk" sampling but with more variance
@@
  • "topk" sampling ("sample" sampling can be more performant but exhibits higher variance)
@@
Finally, we provide more commentary on interesting observations from the grid search results. Structure without Structure Minimum Size enforcing often leads to small functional groups being the most frequent molecular substructures extracted with Beam Enumeration. Enforcing Structure Minimum Size puts it almost on par with Scaffold, suggesting (perhaps not surprisingly) that larger substructures can carry more meaningful information. Moreover, when using "sample" sampling, the generative model undergoes more "filter rounds". Specifically, at each epoch, the sampled batch is filtered to contain the extracted substructures. When using "sample" sampling, the model is more prone to some epochs containing no molecules with the substructures. In practice, this is inconsequential as sampling is computationally inexpensive and a next batch of molecules can easily be sampled. However, specifically in the Structure with "sample" sampling and Structure Minimum Size = 15 experiment, "filter round" can be quite extensive, taking up to 100,000 epochs (maximum observed) for an oracle budget of 5,000 (adding about an hour to the wall time which is minor when the oracle is expensive). This means that many epochs contained molecules without the extracted substructures. There are two observations here: firstly, "sample" sampling can lead to more improbable substructures which are hence less likely to be sampled and secondly, Structure with Structure Minimum Size enforcement leads to extreme biasing (which improves sample efficiency). We believe the remarkable tolerability of the generative model sampling to such bias is an interesting observation. By contrast, Scaffold with Structure Minimum Size enforcement is not as prone to "filter rounds" because Scaffold "truncates" the substructure to its central shape (scaffold). For example, toluene (benzene with a methyl group) has a Bemis-MurckoBemis and Murcko (1996) scaffold of just benzene. The consequence is that Structure leads to more extreme biasing (it is more likely for a molecule to contain benzene than specifically toluene) which is in agreement with the general observation that the diversity of the generated set decreases when using Structure. Overall, both Scaffold and Structure with Structure Minimum Size enforcing exhibits the best performance and "sample" sampling can be more performant than "topk" sampling but exhibits notably higher variance.
@@
All results from the drug discovery case studies are shown in Section D.
B.4 Beam Enumeration: Recommended default Hyperparameters
Taking into consideration all grid search experiments for the illustrative experiment and Drug Discovery case studies, the following optimal hyperparameters are recommended: Patience = 5, Pool Size = 4, Beam Steps = 18, Structure, Structure Minimum Size = 15, "topk" sampling.
Notable differences between the final recommended hyperparameters compared to those found from the illustrative experiment is that Structure and "topk" sampling are more performant than Scaffold and "sample" sampling. In the illustrative experiment, "sample" sampling was sometimes more performant than "topk" sampling. We rationalize these observations as follows: in MPO objectives that include physics-based oracles, structure specificity becomes increasingly important, e.g., specific chemical motifs dock well because they form interactions with the protein. Therefore, "topk" sampling is more robust as there is less variance in the extract substructures compared to "sample" sampling. We empirically observe the increased variance when using "sample" sampling measured by the standard deviation between replicate experiments (Appendix D). In the illustrative experiment where the oracle was more permissive, i.e., any rings saturated with heteroatoms would satisfy the MPO objective, small deviations in the extracted structure do not have as prominent an effect as physics-based oracles which require specificity. Another observation is that Structure sampling often extracts scaffolds with "branch points" which enforces extreme bias that can lead to more focused chemical space exploration. We discuss this in detail in Section D and believe the insights are generally interesting in the context of molecular optimization landscape.
@@ B.4 Beam Enumeration: Recommended default Hyperparameters
B.5 Pseudo-code
The pseudo-code for Beam Enumeration is presented here. The $\oplus$ operator denotes every element on the left is being extended by every element on the right.
@@ B.5 Pseudo-code
Sub-sequences = [Top $k$ <START> Tokens];
@@
Appendix C Illustrative Experiment

### v2

@@ Appendix B Beam Enumeration
B.1 Algorithm Overview
Figure B4: Beam Enumeration overview. a. The proposed method proceeds via 4 steps: 1. generate batch of molecules. 2. filter molecules based on pool to enforce substructure presence, discarding the rest. 3. compute reward 4. update the model. After updating the model, if the reward has improved for consecutive epochs, execute Beam Enumeration. b. Beam Enumeration sequentially enumerates the top $k$ tokens by probability for $N$ beam steps, resulting in an exhaustive set of token sub-sequences. c. All valid substructures (either by the Structure or Scaffold criterion) are extracted from the sub-sequences. The most frequent substructures are used for self-conditioned generation. This overview figure is the same as in the main text.
Beam Enumeration (Fig. B4) is an algorithm that extracts molecular substructures from a generative model’s weights for self-conditioned generation. The problem set-up is any molecular design task to optimize for a target property profile, e.g., high predicted solubility and binding affinity. When molecular generative models are coupled with an optimization algorithm, it should be increasingly likely to generate desirable molecules, i.e., molecules that possess the target property profile.
@@ B.1 Algorithm Overview
In this work, Beam Enumeration is applied to a language-based autoregressive generative model operating on the simplified molecular-input line-entry system (SMILES) (Weininger (1988) representation. The optimization algorithm is Augmented Memory (Guo & Schwaller (2023) which builds on REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020) and casts the optimization process as an on-policy reinforcement learning (RL) problem. Following RL terminology, sampling from the generative model involves sampling trajectories, which in this case, are SMILES, and the desirability of the corresponding molecule is given by the reward.
The underlying hypothesis of Beam Enumeration is that during the RL optimization process, partial trajectories provide a source of signal that can be exploited. Usually, full trajectories are sampled which map to a complete SMILES sequence that can be translated to a molecule. Our assumption is that partial trajectories (partial SMILES sequence) can be mapped to molecular substructures (a part of the full molecule). This statement is not guaranteed as SMILES and molecules are discrete and small perturbations often leads to invalid SMILES. We prove this assumption in Section D by showing that although the vast number of partial trajectories do not map to valid SMILES, the raw number is sufficient to extract a meaningful signal. Correspondingly, Beam Enumeration leverages partial trajectories on the assumption that molecular substructures are on track to becoming full molecules that would receive high reward.
B.2 Enumerating Partial Trajectories
@@
B.3 Extracting Molecular Substructures
Given a set of token sub-sequences, the goal is to extract out the most frequent molecular substructures. This is done by taking each sequence, considering every (sub)-sub-sequence, and counting the number of valid substructures (Fig. B). For example, given the sub-sequence “ABCD”, the set of (sub)-sub-sequences are: “A”, “AB”, “ABC”, and “ABCD”. In practice, we only consider (sub)-sub-sequences with at least three characters (“ABC” and “ABCD”) since each character loosely maps to one atom and three is approximately the minimum for meaningful functional groups, e.g., “C=O”, a carbonyl. The set of most frequent substructures is assumed to be on track to receive a high reward.
B.4 Defining Molecular Substructures: Scaffold vs. Structure
As shown in Fig. B, molecular substructures can be defined on the Scaffold or Structure level. The former extracts the Bemis-Murcko (Bemis & Murcko (1996) scaffold while the latter extracts any valid structure. The any valid structure is an important distinction as our experiments find that extracting by Structure leads to the most frequent molecular substructures being small functional groups that do not have corresponding scaffolds. By contrast, extracting the scaffold always leads to ring structures. Moreover, extracting specifically the Bemis-Murcko scaffold is important as heavy atoms, e.g., nitrogen, are important for biological activity. Consequently, extracted substructures are also enforced to contain at least one heavy atom as we find that benzene, perhaps unsurprisingly, is commonly the most frequent substructure. See Section C for more details on the differing behavior of ‘Scaffold’ vs. ‘Structure’.
B.5 Self-conditioned Generation
@@
B.6 Probabilistic Explainability
The set of most frequent molecular substructures should be meaningful as otherwise, the model’s weights would not have been updated such that these substructures have become increasingly likely to be generated. We verify this statement in the illustrative experiment in the main text and in Section D. In the drug discovery case studies (Appendix E), the extracted substructures are more subtle in why they satisfy the target objective but certainly must possess meaning, however subtle, as otherwise, they would not receive a high reward. In the main text, we show that extracted substructures form core scaffolds and structural motifs in the generated molecules that complement the protein binding cavity. Finally, we emphasize that the correctness and usefulness of this explainability deeply depends on the oracle(s) being optimized for. The extracted substructures do not explain why the generated molecules satisfy the target objective. Rather, they explain why the generated molecules satisfy the oracle. The assumption in a generative design task is that optimizing the oracle is a good proxy for the target objective, e.g., generating molecules that dock well increases the likelihood of the molecules being true binders. This observation directly provides additional commentary on why sample efficiency is so important: the ability to directly optimize expensive high-fidelity oracles would inherently enhance the correctness of the extracted substructures.
Appendix C Beam Enumeration: Findings from Hyperparameter Screening
@@
C.1 Beam Enumeration Hyperparameters
@@
Substructure Type. This hyperparameter has two possible values: Scaffold or Structure. Scaffold extracts Bemis-Murcko (Bemis & Murcko (1996) scaffolds while Structure extracts any valid substructure.
Structure Structure Minimum Size. This hyperparameter enforces the partial SMILES to contain at least a certain number of characters. In effect, this enforces extracted molecular substructures to be larger than a Structure Minimum Size. From the illustrative experiment in the main text and Section D, Structure extraction often leads to small functional groups being the most frequent in the sub-sequences. By enforcing a minimum structure size, Structure extraction leads to partial structures which may carry more meaning. We find that this hyperparameter greatly impacts sample efficiency and we present all our findings in a later sub-section.
@@
Patience. This hyperparameter controls how many successive reward improvements are required before Beam Enumeration executes and molecular substructures are extracted. Recalling the first fact in which Beam Enumeration was proposed on: On a successful optimization trajectory, the model’s weights must change such that desirable molecules are more likely to be generated, on average. Patience is effectively an answer to ”when would extracted substructures be meaningful?” Too low a patience and stochasticity can lead to negative effects while too high a patience diminishes the benefits of Beam Enumeration on sample efficiency.
Token Sampling Method. This hyperparameter has two possible values: ”topk” or ”sample” and denotes how tokens sub-sequences are enumerated. ”topk” takes the top k most probable tokens at each Beam Step while ”sample” samples from the distribution just like during batch generation. Our results show interesting observations surrounding this hyperparameter as ”sample” can work just as well and sometimes even better than taking the ”topk”. These results were unexpected as the underlying hypothesis is that the most probable sub-sequences lead to the most useful substructures being extracted. However, our findings are not in contradiction as sampling the conditional probability distributions would still lead to sampling the top k tokens, on average. Moreover, after extensive experiments, we find that ”sample” leads to more variance in performance across replicates which is in agreement with the assumption that sampling the distributions can lead to more improbable structures. We thoroughly discuss our findings in a later sub-section where we provide hyperparameter recommendations and analyses to the effects of tuning each hyperparameter.
C.2 Hyperparameters: Grid Search
@@
Figure C5: illustrative experiment Generative Yield >0.8. The IntDiv1 (Polykovskiy et al. (2020b) is annotated.
Figure C6: illustrative experiment Generative Yield >0.8 with Structure Minimum Size. The IntDiv1 (Polykovskiy et al. (2020b) is annotated.
Figure C7: illustrative experiment Generative Yield >0.8 with Structure Minimum Size and ”Sample” token sampling. The IntDiv1 (Polykovskiy et al. (2020b) is annotated.
Figure C8: illustrative experiment Oracle Burden (10) >0.8
Figure C9: illustrative experiment Oracle Burden (10) >0.8 with Structure Minimum Size
Figure C10: illustrative experiment Oracle Burden (10) >0.8 with Structure Minimum Size and ”Sample” token sampling
C.3 Analysis of Grid Search Results
@@
  • Scaffold and Structure with Structure Minimum Size: ”sample” sampling can be better than ”topk” sampling but with more variance
@@
  • ”topk” sampling (”sample” sampling can be more performant but exhibits higher variance)
@@
Finally, we provide more commentary on interesting observations from the grid search results. Structure without Structure Minimum Size enforcing often leads to small functional groups being the most frequent molecular substructures extracted with Beam Enumeration. Enforcing Structure Minimum Size puts it almost on par with Scaffold, suggesting (perhaps not surprisingly) that larger substructures can carry more meaningful information. Moreover, when using ”sample” sampling, the generative model undergoes more ”filter rounds”. Specifically, at each epoch, the sampled batch is filtered to contain the extracted substructures. When using ”sample” sampling, the model is more prone to some epochs containing no molecules with the substructures. In practice, this is inconsequential as sampling is computationally inexpensive and a next batch of molecules can easily be sampled. However, specifically in the Structure with ”sample” sampling and Structure Minimum Size = 15 experiment, ”filter round” can be quite extensive, taking up to 100,000 epochs (maximum observed) for an oracle budget of 5,000 (adding about an hour to the wall time which is minor when the oracle is expensive). This means that many epochs contained molecules without the extracted substructures. There are two observations here: firstly, ”sample” sampling can lead to more improbable substructures which are hence less likely to be sampled and secondly, Structure with Structure Minimum Size enforcement leads to extreme biasing (which improves sample efficiency). We believe the remarkable tolerability of the generative model sampling to such bias is an interesting observation. By contrast, Scaffold with Structure Minimum Size enforcement is not as prone to ”filter rounds” because Scaffold ”truncates” the substructure to its central shape (scaffold). For example, toluene (benzene with a methyl group) has a Bemis-Murcko (Bemis & Murcko (1996) scaffold of just benzene. The consequence is that Structure leads to more extreme biasing (it is more likely for a molecule to contain benzene than specifically toluene) which is in agreement with the general observation that the diversity of the generated set decreases when using Structure. Overall, both Scaffold and Structure with Structure Minimum Size enforcing exhibits the best performance and ”sample” sampling can be more performant than ”topk” sampling but exhibits notably higher variance.
@@
All results from the drug discovery case studies are shown in Section E.
C.4 Beam Enumeration: Recommended default Hyperparameters
Taking into consideration all grid search experiments for the illustrative experiment and Drug Discovery case studies, the following optimal hyperparameters are recommended: Patience = 5, Pool Size = 4, Beam Steps = 18, Structure, Structure Minimum Size = 15, ”topk” sampling.
Notable differences between the final recommended hyperparameters compared to those found from the illustrative experiment is that Structure and ”topk” sampling are more performant than Scaffold and ”sample” sampling. In the illustrative experiment, ”sample” sampling was sometimes more performant than ”topk” sampling. We rationalize these observations as follows: in MPO objectives that include physics-based oracles, structure specificity becomes increasingly important, e.g., specific chemical motifs dock well because they form interactions with the protein. Therefore, ”topk” sampling is more robust as there is less variance in the extract substructures compared to ”sample” sampling. We empirically observe the increased variance when using ”sample” sampling measured by the standard deviation between replicate experiments (Appendix E). In the illustrative experiment where the oracle was more permissive, i.e., any rings saturated with heteroatoms would satisfy the MPO objective, small deviations in the extracted structure do not have as prominent an effect as physics-based oracles which require specificity. Another observation is that Structure sampling often extracts scaffolds with ”branch points” which enforces extreme bias that can lead to more focused chemical space exploration. We discuss this in detail in Section E and believe the insights are generally interesting in the context of molecular optimization landscape.
@@ C.4 Beam Enumeration: Recommended default Hyperparameters
C.5 Pseudo-code
This section contains the pseudo-code for Beam Enumeration and Augmented Memory with Beam Enumeration to show the integration.
C.5.1 Beam Enumeration
This subsection contains the Beam Enumeration pseudo-code. The $\oplus$ operator denotes every element on the left is being extended by every element on the right.
@@ C.5.1 Beam Enumeration
Sub-sequences = [Top $k$START¿ Tokens];
@@
C.5.2 Beam Enumeration with Augmented Memory
This subsection contains the Augmented Memory pseudo-code (from Guo et al. (Guo & Schwaller (2023)) with Beam Enumeration integrated. All Beam Enumeration operations are bolded.
[…]
Initialization:
[…]
[3pt] Diversity Filter $DF$;
[3pt] Replay Buffer $B=\{\}$;
[3pt] Substructure Pool $Pool=\{\}$;
[3pt] for $i\leftarrow 1$ to $N$ do
[…]
             Filter sampled batch of SMILES to contain pooled substructures $X_{filtered}$
      [3pt] Compute reward using the scoring function $S(X_{filtered})$;
      [3pt] Modify reward based on the diversity filter $DF(S(X_{filtered}))$;
      [3pt] Update replay buffer $B_{i}=TopK(X_{filtered_{i}}\cup B_{i-1})$;
      [3pt] if reward has improved for $P$ successive epochs then
             Execute Beam Enumeration to update the $Pool$
      [3pt] (Optionally) purge replay buffer;
[…]
Algorithm 2 Beam Enumeration Integrated with Augmented Memory
Appendix D Illustrative Experiment